# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v1

@@ 5.2 Directional Convergence Near Saddle Points
$\mathcal{L}\left(\mathbf{w}_{n},\mathbf{w}_{z}\right)=\frac{1}{2}\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{X};\mathbf{w}_{z})-\mathbf{y}\right\|^{2}.$ (13)
We make the following assumption on the saddle point of (13) considered in this section.
Assumption 2.
$\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$ is a saddle point of (13) such that
$\|\overline{\mathbf{w}}_{n}\|_{2}\in[m,M],\text{ and }\|\overline{\mathbf{w}}_{z}\|_{2}=0,$ (14)
where $m<M$, and $m,M$ are positive constants. Furthermore, there exists $\gamma>0$ such that
$\left\langle\mathbf{w}_{n}-\overline{\mathbf{w}}_{n},\nabla_{\mathbf{w}_{n}}\mathcal{L}(\mathbf{w}_{n},\mathbf{0})\right\rangle\geq 0,\text{ if }\|\mathbf{w}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\leq\gamma.$ (15)
The partition of weights and the above assumption is particularly motivated by aforementioned observation that at saddle points encountered during training, only a certain number of weights are non-zero while others are zero. Clearly, $\overline{\mathbf{w}}_{n}(\overline{\mathbf{w}}_{z})$ represents the set of weights that are non-zero (zero). Additionally, the inequality in (15) is a technical condition to ensure that the gradient flow remains near the saddle point for a sufficiently long time if initialized near the saddle point. We will explain the motivation behind this condition after discussing our main result of this section. The following theorem describes the gradient flow dynamics of (13) when initialized in the neighborhood of a saddle point.
Theorem 5.3.
Let $\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$ satisfy 2 and let $\overline{\mathbf{y}}={\mathbf{y}}-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n}).$ Suppose $\mathbf{p}_{n},\mathbf{p}_{z}$ are fixed unit norm vectors. For any $\epsilon\in(0,\eta/2)$, where $\eta$ is a positive constant, there exist $T_{\epsilon},\tilde{B}_{\epsilon}$, $\tilde{C}_{\epsilon}$ and $\overline{\delta}>0$ such that the following holds: for any $\delta\in(0,\overline{\delta})$ and gradient flow solution $\{\mathbf{w}_{n}(t),\mathbf{w}_{z}(t)\}$ of (13) initialized at $\{\overline{\mathbf{w}}_{n}+\delta\mathbf{p}_{n},\overline{\mathbf{w}}_{z}+\delta\mathbf{p}_{z}\}$, we have
$\|\mathbf{w}_{n}(t)-\overline{\mathbf{w}}_{n}\|_{2}\leq\tilde{C}_{\epsilon}\delta,\text{ and }\|\mathbf{w}_{z}(t)-\overline{\mathbf{w}}_{z}\|_{2}\leq\tilde{B}_{\epsilon}\delta,\text{ for all }t\in\left[0,\frac{T_{\epsilon}}{\delta^{L-2}}\right].$ (16)
Further, for $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$, either
$\|\mathbf{w}_{z}(\overline{T}_{\epsilon})\|_{2}\geq\delta\eta/2,\text{ and }\frac{\mathbf{w}_{z}(\overline{T}_{\epsilon})^{\top}{\mathbf{u}}^{*}}{\|\mathbf{w}_{z}(\overline{T}_{\epsilon})\|_{2}}\geq 1-\left(1+\frac{3}{\eta}\right)\epsilon,$
where ${\mathbf{u}}^{*}$ is a non-negative KKT point of
$\max_{\|\mathbf{u}\|_{2}^{2}=1}\mathcal{N}_{\bar{\mathbf{y}},\mathcal{H}_{z}}(\mathbf{u})=\overline{\mathbf{y}}^{\top}\mathcal{H}_{z}(\mathbf{X};\mathbf{u}),$ (17)
or
$\|\mathbf{w}_{z}(\overline{T}_{\epsilon})\|_{2}\leq 2\delta\epsilon.$
In the above theorem, we choose initialization to be near a saddle point of (13) that satisfies 2, where $\delta>0$ controls the distance of initialization from the saddle point. Since $\overline{\mathbf{w}}_{z}=\mathbf{0}$ and $\mathcal{H}_{z}$ is homogeneous, we have $\mathcal{H}_{z}(\mathbf{x};\overline{\mathbf{w}}_{z})=\mathbf{0}$, Therefore, the vector $\overline{\mathbf{y}}$ represents the residual error at the saddle point $\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$, and plays the same role as $\mathbf{y}$ in Theorem 5.1. In the first part we show that for a given $\epsilon$, we can choose $\delta$ sufficiently small such that gradient flow remains near the saddle point for $t\leq T_{\epsilon}/\delta^{L-2}$. We next describe what happens at $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$. Similar to Theorem 5.1, the weights of small magnitude $\mathbf{w}_{z}$ remains small in norm and either approximately converge in direction to a KKT point of (17), which is the constrained NCF defined with respect to $\overline{\mathbf{y}}$ and $\mathcal{H}_{z}$, or are approximately $\mathbf{0}$. Furthermore, for a fixed training data and neural network, $\mathbf{p}_{z}$ determines which of the above two scenarios will occur, and the positive constant $\eta$ in the above theorem is also determined by $\mathbf{p}_{z}$ and is independent of $\delta$. The proof of the above theorem is similar to proof of Theorem 5.1; see the Appendix C for details. We now turn towards explaining the motivation for (15) in 2. An important step towards establishing directional convergence of $\mathbf{w}_{z}$ is to show that $\mathbf{w}_{n}(t)$ remains near $\overline{\mathbf{w}}_{n}$ and $\mathbf{w}_{z}(t)$ remains small in norm for a sufficiently long time, i.e., (16) holds. Notably, the desired amount of time is inversely proportional to $\delta^{L-2}$. Thus, the closer the initialization is to the saddle point, the longer (16) needs to hold for. Let us assume $\mathbf{w}_{z}(t)$ remains small in norm. In that case, $\mathbf{w}_{n}(t)$ approximately evolves under the gradient flow of $\mathcal{L}(\mathbf{w}_{n},\mathbf{0})$. We use the result of the following lemma to describe the gradient flow dynamics of $\mathcal{L}(\mathbf{w}_{n},\mathbf{0})$ near $\overline{\mathbf{w}}_{n}$.
Lemma 5.4.
Let $\{\overline{\mathbf{w}}_{n},\overline{\mathbf{w}}_{z}\}$ be a saddle point of (13) such that $\|\overline{\mathbf{w}}_{n}\|_{2}\in[m,M],\text{ and }\|\overline{\mathbf{w}}_{z}\|_{2}=0,$ where $m,M$ are positive constants as above. Then, $\overline{\mathbf{w}}_{n}$ is a saddle point of
$\mathcal{L}\left(\mathbf{w}_{n},\mathbf{0}\right)=\frac{1}{2}\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{w}_{n})-\mathbf{y}\right\|^{2}.$ (18)
If $\overline{\mathbf{w}}_{n}$ satisfies (15) for some $\gamma>0$, then $\overline{\mathbf{w}}_{n}$ is a local minimum of (18). Further, if $\mathcal{H}_{n}(.)$ is twice continuously differentiable, and $\overline{\mathbf{w}}_{n}$ is a local minimum of (18) such that $\nabla_{\mathbf{w}_{n}}^{2}\mathcal{L}\left(\overline{\mathbf{w}}_{n},\mathbf{0}\right)\succ\mathbf{0}$, then $\overline{\mathbf{w}}_{n}$ satisfies (15) for some $\gamma>0$.
From the above lemma, we have that $\overline{\mathbf{w}}_{n}$ is a saddle point of $\mathcal{L}(\mathbf{w}_{n},\mathbf{0})$. However, it may not be stable. That is, if $\mathbf{w}_{n}(t)$ is initialized near $\overline{\mathbf{w}}_{n}$, then $\mathbf{w}_{n}(t)$ may not remain near $\overline{\mathbf{w}}_{n}$ for the desired amount of time. However, under (15) $\overline{\mathbf{w}}_{n}$ is a local minimum $\mathcal{L}(\mathbf{w}_{n},\mathbf{0})$, and thus enables stable behavior of gradient flow near $\overline{\mathbf{w}}_{n}$. Hence, (15) ensures that $\mathbf{w}_{n}(t)$ remains near $\overline{\mathbf{w}}_{n}$ for the desired amount of time. We conclude this section by highlighting a major drawback in Theorem 5.3. We have assumed that the weights of the neural network $\mathbf{w}$ can be partitioned into two sets such that the output of the neural network satisfies $\mathcal{H}(\mathbf{x};\mathbf{w})=\mathcal{H}_{n}(\mathbf{x};\mathbf{w}_{n})+\mathcal{H}_{z}(\mathbf{x};\mathbf{w}_{z})$. While this assumption is crucial for our analysis, it excludes fully connected deep neural networks, since such a partition is not possible for those networks. Therefore, Theorem 5.3 is limited to neural networks that have a separable structure. A main reason preventing us from studying gradient flow dynamics near saddle points of fully connected deep neural networks is the lack of theoretical results for such cases. So far, saddle-to-saddle dynamics has only been rigorously studied for shallow networks. Therefore, we do not have a precise knowledge about the saddle points encountered when training fully connected networks. We believe some kind of structural assumptions on the saddle points would be necessary to extend Theorem 5.3 to fully connected networks. We defer investigations along these lines to a future work.
@@ 6 Conclusion and Future Directions
In this paper, we studied the gradient flow dynamics resulting from training deep homogeneous neural networks with small initializations, and established the approximate directional convergence of the weights in the early stages of training. For square loss, we also established similar directional convergence near certain saddle points. However, our results here are not applicable for ReLU networks, so one important future direction would be to rigorously establish similar directional convergence for ReLU networks. Using simple experiments, we highlight the directional convergence of weights, as well as a low-rank structure in the weights of neural network during early stages of training. This latter phenomenon is not explained by our work, and is an intriguing area for future exploration. Another interesting future direction would be to study the entire gradient dynamics of deep neural networks, and our work here could potentially aid in such an endeavor. A crucial step for this would be to extend our results at saddle points for more general fully-connected neural networks.
@@ A.1 Kurdyka-Lojasiewicz Inequality for Definable Functions
Lemma A.1.
[31, Lemma 3.6] Let $f$ be a locally Lipschitz function, definable under an o-minimal structure, and having an open domain $\mathbf{D}\subset\{\mathbf{x}|\|\mathbf{x}\|_{2}>1\}$. For any $c,\eta>0$, there exists a $\nu>0$ and a definable desingularizing function $\Psi$ on $[0,\nu)$ such that
@@ A.2 Euler’s Theorem for Homogeneous Functions
Lemma A.2.
([32, Theorem B.2], [31, Lemma C.1]) Let $F:\mathbb{R}^{k}\rightarrow\mathbb{R}$ be locally Lipschitz, differentiable, and $L-$positively homogeneous for some $L>0$. Then,
@@ A.3 Gronwall’s Inequality
where $\mathcal{J}(\mathbf{X};\mathbf{u}):\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ is the Jacobian of the function $\mathcal{H}(\mathbf{X};\mathbf{u})$. Recall that $\mathcal{S}^{k-1}$ denotes the unit sphere on $\mathbb{R}^{k}$, then we define
@@
We will start by establishing some auxiliary lemmata. The following lemma follows from $L$-homogeneity of $\mathcal{N}(\mathbf{u})$ and Lemma A.2.
Lemma B.1.
@@
The first equality follows by differentiating $\tilde{\mathcal{N}}(\mathbf{u})$ with respect to $\mathbf{u}$; using Lemma B.1 we get the second equality. ∎
@@
Lemma B.3.
If a vector $\mathbf{u}^{*}\in\mathbb{R}^{k\times 1}$ is a first-order KKT point of
@@
$\nabla\mathcal{N}(\mathbf{u}^{*})=\mathcal{J}(\mathbf{X};\mathbf{u}^{*})^{\top}\mathbf{z}=\lambda^{*}\mathbf{u}^{*},\|\mathbf{u}^{*}\|_{2}^{2}=1,$ (23)
where $\lambda^{*}\in\mathbb{R}$ is the Lagrange multiplier. Also, $L\mathcal{N}(\mathbf{u}^{*})=\lambda^{*}$, which implies that $\lambda^{*}\geq 0$ for a non-negative KKT point.
@@
Hence, if $\mathbf{u}^{*}$ is a first-order KKT point then $\|\mathbf{u}^{*}\|_{2}^{2}=1$, and for some $\overline{\lambda}$ we have
$\nabla\mathcal{N}(\mathbf{u}^{*})+2\overline{\lambda}\mathbf{u}^{*}=\mathbf{0},$
@@
$\mathcal{J}(\mathbf{X};\mathbf{u}^{*})^{\top}\mathbf{z}+2\overline{\lambda}\mathbf{u}^{*}=\mathbf{0}.$
We get (23) by choosing $\lambda^{*}=-2\overline{\lambda}$. Further, by Lemma B.1, we have
$\lambda^{*}=\lambda^{*}\|\mathbf{u}^{*}\|_{2}^{2}={\mathbf{u}^{*}}^{\top}\nabla\mathcal{N}(\mathbf{u}^{*})=L\mathcal{N}(\mathbf{u}^{*}).$
Lemma B.4.
@@
Using Lemma B.4, we have
@@
We now turn to proving Lemma 5.2.
Proof of Lemma 5.2.
We begin by showing that either the limit of $\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$ exists or $\mathbf{u}(t)$ converges to $\mathbf{0}$. For this, we consider two cases. Case 1: $\mathcal{N}(\mathbf{u}(0))>0$.
In this case, we first show that $\mathbf{u}(t)$ becomes unbounded at some finite time. Suppose for the sake of contradiction $\|\mathbf{u}(t)\|_{2}<\infty$ for all finite $t\geq 0$. Let $\mathcal{N}(\mathbf{u}(0))=\gamma>0,$ thus $\|\mathbf{u}(0)\|_{2}>0$. Using Lemma B.5, for all $t\geq 0$,
@@
Hence, $\tilde{\mathcal{N}}(\mathbf{u}(t))$ (which we recall is equal to $\mathcal{N}(\mathbf{u}(t))/\|\mathbf{u}(t)\|_{2}^{L}$) is defined for all $t\geq 0$. Now, by Lemma B.4, for all $t\geq 0$, we have
@@
Therefore, using Lemma B.2, for all $t\geq 0$,
@@
$\frac{1}{{L}/{2}-1}\left(\frac{1}{\|\mathbf{u}(0)\|_{2}^{L/2-1}}-\frac{1}{\|\mathbf{u}(t)\|_{2}^{L/2-1}}\right)\geq 2L\tilde{\gamma}t,\forall t\geq 0.$
@@
$\|\mathbf{u}(t)\|_{2}^{L/2-1}\geq\frac{\|\mathbf{u}(0)\|_{2}^{L/2-1}}{1-t\tilde{\gamma}L(L-2)\|\mathbf{u}(0)\|_{2}^{L/2-1}},\forall t\geq 0.$
The above inequality indicates that $\|\mathbf{u}(t)\|_{2}$ is not bounded for all finite $t\geq 0$, which leads to a contradiction. Henceforth, we choose $T^{*}$ to denote the time when $\mathbf{u}(t)$ becomes infinity. Formally, we assume that
@@
We now proceed to prove that $\lim_{t\rightarrow T^{*}}\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$ exists by establishing that the length of the curve swept by $\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$, which is defined as
@@
is of finite length. This proof technique is inspired from [31], and a similar technique was also employed in [1]. From (25) and (26), we know $\tilde{\mathcal{N}}(\mathbf{u}(t))$ is an increasing function in $[0,T^{*})$. Further, since $\mathcal{N}(\mathbf{u})=\mathbf{z}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u})\leq\beta\|\mathbf{z}\|_{2}\|\mathbf{u}\|_{2}^{L}$, where the inequality follows from (21), we have that $\tilde{\mathcal{N}}(\mathbf{u})$ is bounded from above. Thus, using monotone convergence theorem, we have $\lim_{t\rightarrow T^{*}}\tilde{\mathcal{N}}(\mathbf{u}(t))$ exists.
@@
Hence, from (29), we have
@@
which implies $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ exists and is equal to $\frac{\mathbf{u}(T)}{\|\mathbf{u}(T)\|_{2}}$. Thus, we may assume $\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(t))>0$, for all $t\in[0,T^{*})$. We define $g(\mathbf{u})=\overline{f}-\tilde{\mathcal{N}}(\mathbf{u})$. Then, since
@@
Hence, from A.1, there exists a $\nu>0$ and a desingularizing function $\Psi(.)$ defined on $[0,\nu)$ such that if $\|\mathbf{u}\|_{2}>1$ and $g(\mathbf{u})<\nu$, then
$1\geq\Psi^{\prime}(g(\mathbf{u}))\|\mathbf{u}\|_{2}\|{\nabla}g(\mathbf{u})\|_{2}=\Psi^{\prime}(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}))\|\mathbf{u}\|_{2}\|{\nabla}\tilde{\mathcal{N}}(\mathbf{u})\|_{2}.$ (30)
@@
$=\|\mathbf{u}\|_{2}\left\|{\nabla}\tilde{\mathcal{N}}(\mathbf{u})\right\|_{2}\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}\leq\frac{1}{\Psi^{\prime}(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}))}\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}.$
In the above chain of equalities and inequalities, we used Lemma B.2 in the third equality, and the inequality follows from (30). Now, since $\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(t))\in[0,\nu),$ for all $t\in[T,T^{*})$, and $\Psi^{\prime}>0$ on $[0,\nu)$, we can take $\Psi^{\prime}(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(t)))$ to the left hand side to get

### v2

@@ 5.4 Proof Overview
$\mathbf{0}\in$ $\lambda\overline{{\mathbf{W}}}_{1}+\partial_{{\mathbf{W}}_{1}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right)=\lambda{\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}+\partial_{{\mathbf{W}}_{1}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right).$ (16)
Since $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ have non-negative entries, ReLU is positively homogeneous, $\|{\mathbf{a}}_{l}\|_{2}^{2}=1/\sqrt{L}$, and ${\mathbf{a}}_{l}={\mathbf{b}}_{l+1}$, for all $l\in[L-2]$, $q{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, we can write
$\mathcal{H}({\mathbf{x}};{{\mathbf{W}}}_{1},\overline{{\mathbf{W}}}_{2}\cdots,\overline{{\mathbf{W}}}_{L})$ $=\overline{{\mathbf{W}}}_{L}\sigma\left(\overline{{\mathbf{W}}}_{L-1}\cdots\overline{{\mathbf{W}}}_{2}\sigma\left({\mathbf{W}}_{1}{\mathbf{x}}\right)\cdots\right)$
$=q{\mathbf{a}}_{L-1}^{\top}\sigma({\mathbf{a}}_{L-1}{\mathbf{a}}_{L-2}^{\top}\sigma({\mathbf{a}}_{L-2}{\mathbf{a}}_{L-3}^{\top}\cdots\sigma({\mathbf{a}}_{2}{\mathbf{a}}_{1}^{\top}\sigma({\mathbf{W}}_{1}{\mathbf{x}}))\cdots)/L^{1/4}$
$=q\|{\mathbf{a}}_{L-1}\|_{2}^{2}\|{\mathbf{a}}_{L-2}\|_{2}^{2}\cdots\|{\mathbf{a}}_{2}\|_{2}^{2}\sigma^{L-2}({\mathbf{a}}_{1}^{\top}\sigma({\mathbf{W}}_{1}{\mathbf{x}}))/L^{1/4}$
$=q\left({1}/{\sqrt{L}}\right)^{L-2}\sigma^{L-2}({\mathbf{a}}_{1}^{\top}\sigma({\mathbf{W}}_{1}{\mathbf{x}}))/L^{1/4}.$
We use the above equality to simplify $\partial_{{\mathbf{W}}_{1}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right)$ and use it in eq. 16 to prove ${\mathbf{b}}_{1}$ is a KKT point of the smaller optimization problem. Also, note that since ReLU activation is non-differentiable, we used the Clarke sub-differential in the KKT conditions. A challenging aspect of using Clarke sub-differentials is that the chain rule does not hold with equality, instead it yields a set and the true sub-differential belongs to that set (see Lemma 15). Consequently, it is not feasible to simplify $\partial_{{\mathbf{W}}_{1}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right)$ using the above equality by simply applying the chain rule. We follow a more direct approach to simplify the Clarke sub-differential, which uses positive homogeneity of the activation function and non-negativity of ${\mathbf{a}}_{1}$.
@@ 6 Conclusion and Future Directions
In this paper, we studied the gradient flow dynamics resulting from training deep homogeneous neural networks with small initializations, and established the approximate directional convergence of the weights in the early stages of training. However, these results are not applicable for ReLU networks, so one important future direction would be to rigorously establish similar directional convergence for ReLU networks.
We also derived necessary and sufficient conditions for the KKT points of the constrained NCF that have rank-one hidden weights, for feed-forward neural networks with Leaky ReLU and polynomial Leaky ReLU activation functions. The constrained NCF is a non-convex problem and all KKT points may not be low-rank; however, our experiments suggest that gradient ascent tends to converge towards low-rank KKT points. Understanding this phenomenon is also an important area for future exploration, which will also help us understand why gradient descent on training loss converges towards low-rank KKT points of the NCF in the early stages of training. We hope our characterization of rank-one KKT points will aid in this exploration.
@@ A.1 Kurdyka-Lojasiewicz inequality for definable functions
Lemma 6.
(Ji & Telgarsky, 2020, Lemma 3.6) Let $f$ be a locally Lipschitz function, definable under an o-minimal structure, and having an open domain $\mathbf{D}\subset\{\mathbf{x}|\|\mathbf{x}\|_{2}>1\}$. For any $c,\eta>0$, there exists a $\nu>0$ and a definable desingularizing function $\Psi$ on $[0,\nu)$ such that
@@ A.2 Euler’s Theorem for homogeneous functions
Lemma 7.
((Lyu & Li, 2020, Theorem B.2), (Ji & Telgarsky, 2020, Lemma C.1)) Let $F:\mathbb{R}^{k}\rightarrow\mathbb{R}$ be locally Lipschitz, differentiable, and $L-$positively homogeneous for some $L>0$. Then,
@@ B.1 Proof of Lemma 2
where $\mathcal{J}(\mathbf{X};\mathbf{u}):\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ is the Jacobian of the function $\mathcal{H}(\mathbf{X};\mathbf{u})$. Let $\mathcal{S}^{k-1}$ denote the unit sphere on $\mathbb{R}^{k}$, and define
@@
We will start by establishing some auxiliary lemmata. The following lemma follows from $L$-homogeneity of $\mathcal{N}(\mathbf{u})$ and Lemma 7.
Lemma 9.
@@
The first equality follows by differentiating $\tilde{\mathcal{N}}(\mathbf{u})$ with respect to $\mathbf{u}$; using Lemma 9 we get the second equality. ∎
@@
Lemma 11.
If a vector $\mathbf{u}_{*}\in\mathbb{R}^{k\times 1}$ is a first-order KKT point of
@@
$\nabla\mathcal{N}(\mathbf{u}_{*})=\mathcal{J}(\mathbf{X};\mathbf{u}_{*})^{\top}\mathbf{z}=\lambda^{*}\mathbf{u}_{*},\|\mathbf{u}_{*}\|_{2}^{2}=1,$ (21)
where $\lambda^{*}\in\mathbb{R}$ is the Lagrange multiplier. Also, $L\mathcal{N}(\mathbf{u}_{*})=\lambda^{*}$, which implies that $\lambda^{*}\geq 0$ for a non-negative KKT point.
@@
Hence, if $\mathbf{u}_{*}$ is a first-order KKT point then $\|\mathbf{u}_{*}\|_{2}^{2}=1$, and for some $\overline{\lambda}$ we have
$\nabla\mathcal{N}(\mathbf{u}_{*})+2\overline{\lambda}\mathbf{u}_{*}=\mathbf{0},$
@@
$\mathcal{J}(\mathbf{X};\mathbf{u}_{*})^{\top}\mathbf{z}+2\overline{\lambda}\mathbf{u}_{*}=\mathbf{0}.$
We get eq. 21 by choosing $\lambda^{*}=-2\overline{\lambda}$. Further, by Lemma 9, we have
$\lambda^{*}=\lambda^{*}\|\mathbf{u}_{*}\|_{2}^{2}={\mathbf{u}_{*}}^{\top}\nabla\mathcal{N}(\mathbf{u}_{*})=L\mathcal{N}(\mathbf{u}_{*}).$
Lemma 12.
@@
Using Lemma 12, we have
@@
We now turn to proving Lemma 2.
Proof of Lemma 2:.
We begin by showing that either the limit of $\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$ exists or $\mathbf{u}(t)$ converges to $\mathbf{0}$. For this, we consider two cases.
Case 1: $\mathcal{N}(\mathbf{u}(0))>0$.
In this case, we first show that $\mathbf{u}(t)$ becomes unbounded at some finite time. Suppose for the sake of contradiction $\|\mathbf{u}(t)\|_{2}<\infty$ for all finite $t\geq 0$.
Let $\mathcal{N}(\mathbf{u}(0))=\gamma>0,$ thus $\|\mathbf{u}(0)\|_{2}>0$. Using Lemma 13, for all $t\geq 0$,
@@
Hence, $\tilde{\mathcal{N}}(\mathbf{u}(t))$ (which we recall is equal to $\mathcal{N}(\mathbf{u}(t))/\|\mathbf{u}(t)\|_{2}^{L}$) is defined for all $t\geq 0$. Now, by Lemma 12, for all $t\geq 0$, we have
@@
Therefore, using Lemma 10, for all $t\geq 0$,
@@
$\frac{1}{{L}/{2}-1}\left(\frac{1}{\|\mathbf{u}(0)\|_{2}^{L-2}}-\frac{1}{\|\mathbf{u}(t)\|_{2}^{L-2}}\right)\geq 2L\tilde{\gamma}t,\forall t\geq 0.$
@@
$\|\mathbf{u}(t)\|_{2}^{L-2}\geq\frac{\|\mathbf{u}(0)\|_{2}^{L-2}}{1-t\tilde{\gamma}L(L-2)\|\mathbf{u}(0)\|_{2}^{L-2}},\forall t\geq 0.$
The above inequality indicates that $\|\mathbf{u}(t)\|_{2}$ is not bounded for all finite $t\geq 0$, which leads to a contradiction.
Henceforth, we choose $T^{*}$ to denote the time when $\mathbf{u}(t)$ becomes infinity. Formally, we assume that
@@
We now proceed to prove that $\lim_{t\rightarrow T^{*}}\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$ exists by establishing that the length of the curve swept by $\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$, given by
@@
is of finite length. This proof technique is inspired from Ji & Telgarsky (2020), and a similar technique was also employed in Kumar & Haupt (2024).
From eq. 23 and eq. 24, we know $\tilde{\mathcal{N}}(\mathbf{u}(t))$ is an increasing function in $[0,T^{*})$. Further, since $\mathcal{N}(\mathbf{u})=\mathbf{z}^{\top}\mathcal{H}(\mathbf{X};\mathbf{u})\leq\beta\|\mathbf{z}\|_{2}\|\mathbf{u}\|_{2}^{L}$, where the inequality follows from eq. 19, we have that $\tilde{\mathcal{N}}(\mathbf{u})$ is bounded from above. Thus, using monotone convergence theorem, we have $\lim_{t\rightarrow T^{*}}\tilde{\mathcal{N}}(\mathbf{u}(t))$ exists.
@@
Hence, from eq. 27, we have
@@
which implies $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ exists and is equal to $\frac{\mathbf{u}(T)}{\|\mathbf{u}(T)\|_{2}}$.
Thus, we may assume $\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(t))>0$, for all $t\in[0,T^{*})$. We define $g(\mathbf{u})=\overline{f}-\tilde{\mathcal{N}}(\mathbf{u})$. Then, since
@@
Hence, from Lemma 6, there exists a $\nu>0$ and a desingularizing function $\Psi(.)$ defined on $[0,\nu)$ such that if $\|\mathbf{u}\|_{2}>1$ and $g(\mathbf{u})<\nu$, then
$1\leq\Psi^{\prime}(g(\mathbf{u}))\|\mathbf{u}\|_{2}\|{\nabla}g(\mathbf{u})\|_{2}=\Psi^{\prime}(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}))\|\mathbf{u}\|_{2}\|{\nabla}\tilde{\mathcal{N}}(\mathbf{u})\|_{2}.$ (28)
@@
$=\|\mathbf{u}\|_{2}\left\|{\nabla}\tilde{\mathcal{N}}(\mathbf{u})\right\|_{2}\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}\geq\frac{1}{\Psi^{\prime}(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}))}\left\|\frac{d}{dt}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}.$
In the above chain of equalities and inequalities, we used Lemma 10 in the third equality, and the inequality follows from eq. 28. Now, since $\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(t))\in[0,\nu),$ for all $t\in[T,T^{*})$, and $\Psi^{\prime}>0$ on $[0,\nu)$, we can take $\Psi^{\prime}(\overline{f}-\tilde{\mathcal{N}}(\mathbf{u}(t)))$ to the left hand side to get